OpenStudy (anonymous):

how would you solve t^2 - 4t - 165 = 0 by completing the square?

OpenStudy (mertsj):

\[t^2-4t+4=165+4\] \[(t-2)^2=169\]

OpenStudy (mertsj):

Now take the square root of both sides and don't forget there are 2 square roots of 169

OpenStudy (callisto):

t^2 - 4t - 165 = 0 (t^2 - 4t +4) -4-165 =0 (t-2)^2 -169 =0 (t-2)^2=169 (t-2) = 13 or (t-2) = -13 t = 15 or t = -11

OpenStudy (anonymous):

thank you

OpenStudy (mertsj):

yw